# The Oracle Complexity of Simplex-based Matrix Games: Linear Separability and Nash Equilibria

### v2

@@ Abstract
We study the problem of solving matrix games of the form $\max_{\mathbf{w}\in\mathcal{W}}\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}A\mathbf{w}$, where $A$ is some matrix and $\Delta$ is the probability simplex. This problem encapsulates canonical tasks such as finding a linear separator and computing Nash equilibria in zero-sum games. However, perhaps surprisingly, its inherent complexity (as formalized in the standard framework of oracle complexity (Nemirovski and Yudin, 1983)) is not well-understood. In this work, we first identify different oracle models which are implicitly used by prior algorithms, amounting to multiplying the matrix $A$ by a vector from either one or both sides. We then prove complexity lower bounds for algorithms under both access models, which in particular imply a separation between them. Specifically, we start by showing that algorithms for linear separability based on one-sided multiplications must require $\Omega(\gamma_{A}^{-2})$ iterations, where $\gamma_{A}$ is the margin, as matched by the Perceptron algorithm. We then prove that accelerated algorithms for this task, which utilize multiplications from both sides, must require $\tilde{\Omega}(\gamma_{A}^{-2/3})$ iterations, establishing the first oracle complexity barrier for such algorithms. Finally, by adapting our lower bound to $\ell_{1}$ geometry, we prove that computing an $\epsilon$-approximate Nash equilibrium requires $\tilde{\Omega}(\epsilon^{-2/5})$ iterations, which is an exponential improvement over the previously best-known lower bound due to Hadiji et al. (2024).
@@ 1 Introduction
  • Nash equilibria with a two-sided oracle (Theorem 5.2): We prove that under the two-sided oracle model, any deterministic algorithm requires $\tilde{\Omega}(\epsilon^{-2/5})$ oracle calls in order to find an $\epsilon$-suboptimal strategy. Hence, the same lower bound holds for finding an $\epsilon$-approximate Nash equilibrium, which is an exponential improvement over the previously known lower bound by Hadiji et al. (2024). The proof is based on an adaptation of the linear separability lower bound technique to an $\ell_{1}$ geometry. Along the way, we also prove an identical lower bound for $\ell_{1}$/simplex matrix games, where $\mathcal{W}$ in Eq. (1) is the unit $\ell_{1}$ ball.
@@ 2 Preliminaries
[…]
@@
As described in the introduction, we study the complexity of matrix games via the well-known optimization-theoretic frameork of oracle complexity (Nemirovski and Yudin, 1983). This framework focuses on iterative methods, where each iteration utilizes some restricted information about the relevant objective function. In our context of matrix games as in Eq. (1), we assume that the domain $\mathcal{W}$ is fixed and known, and that the matrix $A$ is known to belong to some set $\mathscr{A}$: For linear separability, in light of existing methods and upper bounds, it is natural to consider all $n\times d$ matrices whose rows have Euclidean norm at most $1$. Similarly, for Nash equilibrium, it is natural to consider all $n\times d$ matrices whose entries have values in $[-1,+1]$. Crucially, the algorithm has no additional prior knowledge of $A$. In order to solve the matrix problem, the algorithm has access to an oracle $\mathcal{O}(\cdot)$ which provides some limited information about $A$: For example, given the vectors $\mathbf{p}\in\mathbb{R}^{n},\mathbf{w}\in\mathbb{R}^{d}$ chosen by the algorithm, the oracle returns $\mathbf{p}^{\top}A$ and $A\mathbf{w}$. The algorithm interacts with the oracle for a given number of iterations, after which it returns an output as a function of all previous oracle responses. One can then ask how many iterations are required by some algorithm, for the output to satisfy a certain performance metric for all $A\in\mathscr{A}$ (as a function of the problem parameters $n,d,\mathscr{A},\mathcal{W}$ and the type of oracle $\mathcal{O}(\cdot)$). This framework naturally models standard scalable approaches for solving matrix games, and allows one to prove both upper bounds and unconditional lower bounds for iterative algorithms, assuming they interact with the matrix $A$ in a manner corresponding to the given oracle.
@@ 3 Oracle models
Thus, the Perceptron’s convergence guarantee implies that $O(\gamma_{A}^{-2})$ queries to a one-sided oracle is sufficient for finding a linear separator. In fact, a more general result can be achieved by applying the well-known subgradient method on the equivalent convex problem $\min_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\max_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}(-A)\mathbf{w}$. Specifically, by standard results, this method requires $O(1/\epsilon^{2})$ subgradient computations in order to find a vector $\mathbf{w}$ whose value is $\epsilon$-suboptimal (Nesterov, 2018). This holds for any $\epsilon$, whereas the guarantee for the Perceptron is merely for the special case $\epsilon=\gamma_{A}$ (namely, we seek a solution whose value is $>0$, with the optimal value being $\gamma_{A}$). From an oracle-complexity perspective, implementing such methods requires access to supergradients of the function $f(\mathbf{w})=\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}=\min_{l\in[n]}(A\mathbf{w})_{l}$, which equal $A_{l_{\min}}$ where $l_{\min}\in\arg\min_{l\in[n]}(A\mathbf{w})_{l}$. Thus, we can model these methods as iteratively interacting with the following supergradient oracle:
@@ 4.1 Oracle complexity with a one-sided oracle
We note that since $2T+1<d$, the dimensionality is sufficiently high to find orthogonal unit vectors as specified. Moreover, it is easy to verify that because of the orthogonality, it holds that $A_{t}\mathbf{w}_{t}=A\mathbf{w}_{t}$ for all $t\in[T]$: Namely, the responses given to the algorithm are consistent with the oracle responses on the matrix $A$. However, after $T$ iterations, $|\mathcal{L}_{T}|\leq T$, yet there are $T+1$ rows. Therefore, at least one row of $A$ will be chosen to be orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{T+1}$, and in particular to $\mathbf{w}_{T+1}$. Hence, $A\mathbf{w}_{T+1}$ contains a 00 entry, so
@@ 4.2 Oracle complexity with a two-sided oracle
$\max_{l\in[n]}\|A_{l}\|\leq 1~{}~{}~{}\text{and}~{}~{}~{}\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{1}{cT\sqrt{T\log(n)}}~{},$
@@
As before, we can state this lower bounds in terms of a margin parameter $\gamma_{A}$: Given any small enough $\gamma_{A}$, we can choose $T$ such that $1/(cT\sqrt{T\log(n)})\geq\gamma_{A}$, and get that for any algorithm, there exists an $n\times d$ matrix $A$ (for large enough $n,d$) with margin parameter at least $\gamma_{A}$ and rows of norm at most $1$, such that the required number of iterations $T$ to find a linear separator is $\Omega(\log^{-1/3}(n)\cdot\gamma_{A}^{-2/3})$.
@@
Note that by construction, $\mathbf{v}_{t}$ is a standard Gaussian random vector in $\mathbb{R}^{n}$, scaled by $\beta$ and projected on the subspace orthogonal to $\mathbf{v}_{0},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t}$ (hence, $\mathbf{v}_{t}$ is orthogonal to all these vectors). This is possible by the assumption $2T+1\leq n$, which implies that $M_{t}$ is always a “thin” matrix representing the basis of a strict subspace of $\mathbb{R}^{n}$. Similarly, the assumption $2T+1\leq d$ implies that choosing $\mathbf{u}_{t}\in\mathbb{R}^{d}$ as orthogonal to $12t-$ 1 given vectors (as described above) is indeed possible. The choice of a Gaussian distribution is crucially used in order to control the sign and magnitudes of various quantities associated with the matrix, as will be seen later in the proof.
@@ 5 Nash equilibria: simplex/simplex games
As previously discussed, the objective above corresponds to finding a Nash equilibrium in zero-sum bilinear games. In this section we will assume that each entry of $A$ is in $[-1,+1]$, which is the standard normalization for this setting (cf. Nemirovski, 2004; Hadiji et al., 2024).
@@
The formal proof appears in the appendix. In a nutshell, the idea is that given a matrix $A\in[-1,+1]^{n\times d}$ and an algorithm for a simplex/simplex matrix game, we can feed the algorithm with the matrix $(A;-A)\in\mathbb{R}^{n\times 2d}$, and convert the algorithm’s output $\mathbf{w}$ (a vector in the $(2d-1)$-simplex) to a $d$-dimensional vector $\mathbf{w}^{\prime}$ in the $\ell_{1}$ unit ball, so that $A\mathbf{w}^{\prime}=(A;-A)\mathbf{w}$, leading to similar guarantees. We note that the reduction does require us to modify the matrix width $d$ by a factor of $2$, but this will not affect our lower bound by more than a small constant factor (as the bound we will show applies to any sufficiently large $d$). We also note that the computational complexity of the two algorithms in the reduction are essentially identical.
@@
The following holds for some large enough universal constant $c>0$: For any $T>c$, suppose $d,n$ are sufficiently large so that $d>cT$ and $n>cT^{2}\log(T)$. Then for any deterministic algorithm for solving Eq. (4), there exists a matrix $A\in[-1,+1]^{n\times d}$ satisfying
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{1}{c\log(d)\sqrt{\log(n)}}\cdot\frac{1}{T^{2}\sqrt{T}}~{},$
@@
In particular, by fixing some small enough $\epsilon>0$ and choosing $T$ such that the lower bound in the theorem is at least $\epsilon$, we can restate the bound as follows: If the number of iterations $T$ is less than $\Omega(\log(d)^{-2/5}\log(n)^{-1/5}\epsilon^{-2/5})$, there exists a matrix $A$ such that the optimal value is at least $\epsilon$, yet the algorithm’s output has value less than 00, hence is suboptimal by at least $\epsilon$. As discussed previously, by Proposition 5.1, this lower bound for $\ell_{1}$/simplex automatically extends to simplex/simplex games, up to a small numerical constant which is absorbed in the $\Omega(\cdot)$ notation.
To prove the theorem, we utilize a construction very similar to the $\ell_{2}$ case in the previous section, except that the vectors need to be scaled differently, to satisfy the different constraints on $A$ and $\mathbf{w}$. Moreover, since now the constraint on $A$ is that each individual entry is in $[-1,+1]$, we need to choose each $\mathbf{u}_{t}$ in Construction 1 more carefully, as follows:
@@
By construction, $\mathbf{u}_{t}$ still satisfies the desideratum that it is a unit vector orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{v}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$. However, this particular construction gives us probabilistic control over the magnitude of its individual entries. Specifically, we have the following:
@@
For $A$ as defined in Construction 2, and for any $\delta\in(0,1)$, it holds with probability at least $1-2T\left(\delta+\exp(-d/48)\right)$ with respect to the construction that
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~{}\geq~{}\frac{\alpha}{2T\sqrt{8d\log(2d/\delta)}}~{}.$
@@
Suppose $A$ is constructed as in Construction 2. Then for any $\delta\in(0,1)$ such that
$\sqrt{\frac{8\log(2d/\delta)}{d}}\cdot\left(\alpha+2T\beta\sqrt{2\log(2Tn/\delta)}\right)~{}\leq~{}1~{},$
it holds with probability at least $1-3T\left(\delta+\exp(-d/48)\right)$ that $A\in[-1,+1]^{n\times d}$.

### v3

@@ Abstract
We study the problem of solving matrix games of the form $\max_{\mathbf{w}\in\mathcal{W}}\min_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}A\mathbf{w}$, where $A$ is some matrix and $\Delta$ is the probability simplex. This problem encapsulates canonical tasks such as finding a linear separator and computing Nash equilibria in zero-sum games. However, perhaps surprisingly, its inherent complexity (as formalized in the standard framework of oracle complexity (Nemirovski and Yudin, 1983)) is not well-understood. In this work, we first identify different oracle models which are implicitly used by prior algorithms, amounting to multiplying the matrix $A$ by a vector from either one or both sides. We then prove complexity lower bounds for algorithms under both access models, which in particular imply a separation between them. Specifically, we start by showing that algorithms for linear separability based on one-sided multiplications must require $\Omega(\gamma_{A}^{-2})$ iterations, where $\gamma_{A}$ is the margin, as matched by the Perceptron algorithm. We then prove that accelerated algorithms for this task, which utilize multiplications from both sides, must require $\tilde{\Omega}(\gamma_{A}^{-2/3})$ iterations, establishing the first oracle complexity barrier for such algorithms. Finally, by adapting our lower bound to $\ell_{1}$ geometry, we prove that computing an $\epsilon$-approximate Nash equilibrium requires $\tilde{\Omega}(\epsilon^{-2/3})$ iterations, which is an exponential improvement over the previously best-known lower bound due to Hadiji et al. (2024).
@@ 1 Introduction
  • Nash equilibria with a two-sided oracle (Theorem 5.2): We prove that under the two-sided oracle model, any deterministic algorithm requires $\tilde{\Omega}(\epsilon^{-2/3})$ oracle calls in order to find an $\epsilon$-suboptimal strategy. Hence, the same lower bound holds for finding an $\epsilon$-approximate Nash equilibrium, which is an exponential improvement over the previously known lower bound by Hadiji et al. (2024). The proof is based on an adaptation of the linear separability lower bound technique to an $\ell_{1}$ geometry. Along the way, we also prove an identical lower bound for $\ell_{1}$/simplex matrix games, where $\mathcal{W}$ in Eq. (1) is the unit $\ell_{1}$ ball.
@@ 2 Preliminaries
We use capital letters to denote matrices, and bold-face letters to denote vectors. Vectors are always in column form. Given an indexed vector $\mathbf{v}_{t}$, $v_{t,i}$ denotes its $i$-th entry. $\mathbf{e}_{i}$ denotes the $i$-th standard basis vector. $\mathbf{1}$ is the all-ones vector. Given a matrix $A$, $A_{l}$ refers to its $l$-th row. $\|\cdot\|$ refers to the Euclidean norm $\|\cdot\|_{2}$, $\|\cdot\|_{p}$ refers to the $\ell_{p}$ norm, and $\log(\cdot)$ refers to the natural logarithm, unless specified otherwise. Finally, $[n]$ is shorthand for $\{1,\ldots,n\}$.
@@
As described in the introduction, we study the complexity of matrix games via the well-known optimization-theoretic frameork of oracle complexity (Nemirovski and Yudin, 1983). This framework focuses on iterative methods, where each iteration utilizes some restricted information about the relevant objective function. In our context of matrix games as in Eq. (1), we assume that the domain $\mathcal{W}$ is fixed and known, and that the matrix $A$ is known to belong to some set $\mathscr{A}$: For linear separability, in light of existing methods and upper bounds, it is natural to consider all $n\times d$ matrices whose rows have Euclidean norm at most $1$. Similarly, for Nash equilibrium, it is natural to consider all $n\times d$ matrices whose entries have values in [1,+1][-1,+1]. Crucially, the algorithm has no additional prior knowledge of $A$. In order to solve the matrix problem, the algorithm has access to an oracle $\mathcal{O}(\cdot)$ which provides some limited information about $A$: For example, given the vectors $\mathbf{p}\in\mathbb{R}^{n},\mathbf{w}\in\mathbb{R}^{d}$ chosen by the algorithm, the oracle returns $\mathbf{p}^{\top}A$ and $A\mathbf{w}$. The algorithm interacts with the oracle for a given number of iterations, after which it returns an output as a function of all previous oracle responses. One can then ask how many iterations are required by some algorithm, for the output to satisfy a certain performance metric for all $A\in\mathscr{A}$ (as a function of the problem parameters $n,d,\mathscr{A},\mathcal{W}$ and the type of oracle $\mathcal{O}(\cdot)$). This framework naturally models standard scalable approaches for solving matrix games, and allows one to prove both upper bounds and unconditional lower bounds for iterative algorithms, assuming they interact with the matrix $A$ in a manner corresponding to the given oracle.
@@ 3 Oracle models
Thus, the Perceptron’s convergence guarantee implies that $O(\gamma_{A}^{-2})$ queries to a one-sided oracle is sufficient for finding a linear separator. In fact, a more general result can be achieved by applying the well-known subgradient method on the equivalent convex problem $\min_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\max_{\mathbf{p}\in\Delta}\mathbf{p}^{\top}(-A)\mathbf{w}$. Specifically, by standard results, this method requires $O(1/\epsilon^{2})$ subgradient computations in order to find a vector $\mathbf{w}$ whose value is $\epsilon$-suboptimal (Nesterov, 2018). This holds for any $\epsilon$, whereas the guarantee for the Perceptron is merely for the special case $\epsilon=\gamma_{A}$ (namely, we seek a solution whose value is >0>0, with the optimal value being $\gamma_{A}$). From an oracle-complexity perspective, implementing such methods requires access to supergradients of the function $f(\mathbf{w})=\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}=\min_{l\in[n]}(A\mathbf{w})_{l}$, which equal $A_{l_{\min}}$ where $l_{\min}\in\arg\min_{l\in[n]}(A\mathbf{w})_{l}$. Thus, we can model these methods as iteratively interacting with the following supergradient oracle:
@@ 4.1 Oracle complexity with a one-sided oracle
We note that since $2T+1<d$, the dimensionality is sufficiently high to find orthogonal unit vectors as specified. Moreover, it is easy to verify that because of the orthogonality, it holds that $A_{t}\mathbf{w}_{t}=A\mathbf{w}_{t}$ for all $t\in[T]$: Namely, the responses given to the algorithm are consistent with the oracle responses on the matrix $A$. However, after $T$ iterations, $|\mathcal{L}_{T}|\leq T$, yet there are $T+1$ rows. Therefore, at least one row of $A$ will be chosen to be orthogonal to $\mathbf{w}_{1},\ldots,\mathbf{w}_{T+1}$, and in particular to $\mathbf{w}_{T+1}$. Hence, $A\mathbf{w}_{T+1}$ contains a 0 entry, so
@@ 4.2 Oracle complexity with a two-sided oracle
$\max_{l\in[n]}\|A_{l}\|\leq 1~~~\text{and}~~~\max_{\mathbf{w}:\|\mathbf{w}\|\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{cT^{3/2}\sqrt{\log(n)}}~,$
@@
As before, we can state this lower bounds in terms of a margin parameter $\gamma_{A}$: Given any small enough $\gamma_{A}$, we can choose $T$ such that $1/(cT^{3/2}\sqrt{\log(n)})\geq\gamma_{A}$, and get that for any algorithm, there exists an $n\times d$ matrix $A$ (for large enough $n,d$) with margin parameter at least $\gamma_{A}$ and rows of norm at most $1$, such that the required number of iterations $T$ to find a linear separator is $\Omega(\log^{-1/3}(n)\cdot\gamma_{A}^{-2/3})$.
@@
Note that by construction, $\mathbf{v}_{t}$ is a standard Gaussian random vector in $\mathbb{R}^{n}$, scaled by $\beta$ and projected on the subspace orthogonal to $\mathbf{v}_{0},\ldots,\mathbf{v}_{t-1},\mathbf{p}_{1},\ldots,\mathbf{p}_{t}$ (hence, $\mathbf{v}_{t}$ is orthogonal to all these vectors). This is possible by the assumption $2T+1\leq n$, which implies that $M_{t}$ is always a “thin” matrix representing the basis of a strict subspace of $\mathbb{R}^{n}$. Similarly, the assumption $2T+1\leq d$ implies that choosing $\mathbf{u}_{t}\in\mathbb{R}^{d}$ as orthogonal to 2t12t-1 given vectors (as described above) is indeed possible. The choice of a Gaussian distribution is crucially used in order to control the sign and magnitudes of various quantities associated with the matrix, as will be seen later in the proof.
@@ 5 Nash equilibria: simplex/simplex games
As previously discussed, the objective above corresponds to finding a Nash equilibrium in zero-sum bilinear games. In this section we will assume that each entry of $A$ is in [1,+1][-1,+1], which is the standard normalization for this setting (cf. Nemirovski, 2004; Hadiji et al., 2024).
@@
The formal proof appears in the appendix. In a nutshell, the idea is that given a matrix $A\in[-1,+1]^{n\times d}$ and an algorithm for a simplex/simplex matrix game, we can feed the algorithm with the matrix $(A;-A)\in\mathbb{R}^{n\times 2d}$, and convert the algorithm’s output $\mathbf{w}$ (a vector in the (2d1)(2d-1)-simplex) to a $d$-dimensional vector $\mathbf{w}^{\prime}$ in the $\ell_{1}$ unit ball, so that $A\mathbf{w}^{\prime}=(A;-A)\mathbf{w}$, leading to similar guarantees. We note that the reduction does require us to modify the matrix width $d$ by a factor of $2$, but this will not affect our lower bound by more than a small constant factor (as the bound we will show applies to any sufficiently large $d$). We also note that the computational complexity of the two algorithms in the reduction are essentially identical.
@@
The following holds for some large enough universal constant $c>0$: For any $T>c$, suppose $n,d$ are sufficiently large so that $d>cT$ and $n>cT^{2}\log(Tdn)$. Then for any deterministic algorithm for solving Eq. (4), there exists a matrix $A\in[-1,+1]^{n\times d}$ satisfying
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}~\geq~\frac{1}{c\log(dn)}\cdot\frac{1}{T^{3/2}}~,$
@@
In particular, by fixing some small enough $\epsilon>0$ and choosing $T$ such that the lower bound in the theorem is at least $\epsilon$, we can restate the bound as follows: If the number of iterations $T$ is less than $\Omega(\log(dn)^{-2/3}\epsilon^{-2/3})$, there exists a matrix $A$ such that the optimal value is at least $\epsilon$, yet the algorithm’s output has value less than 0, hence is suboptimal by at least $\epsilon$. As discussed previously, by Proposition 5.1, this lower bound for $\ell_{1}$/simplex automatically extends to simplex/simplex games, up to a small numerical constant which is absorbed in the $\Omega(\cdot)$ notation.
To prove the theorem, we utilize a construction very similar to the $\ell_{2}$ case in the previous section, except that the vectors need to be scaled differently, to satisfy the different constraints on $A$ and $\mathbf{w}$. Moreover, since now the constraint on $A$ is that each individual entry is in [1,+1][-1,+1], we need to choose each $\mathbf{u}_{t}$ in Construction 1 more carefully, as follows:
@@
By construction, $\mathbf{u}_{t}$ still satisfies the desideratum that it is a unit vector orthogonal to $\mathbf{u}_{1},\ldots,\mathbf{u}_{t-1}$ and $\mathbf{w}_{1},\ldots,\mathbf{w}_{t}$. However, this particular construction gives us probabilistic control over the magnitude of its individual entries. Specifically, we have the following:
@@
For $A$ as defined in Construction 2, it holds that
$\max_{\mathbf{w}:\|\mathbf{w}\|_{1}\leq 1}\min_{\mathbf{p}\in\Delta^{n-1}}\mathbf{p}^{\top}A\mathbf{w}\geq\frac{\alpha}{\sqrt{d(T+1)}}~.$
@@
Suppose $A$ is constructed as in Construction 2. Then there is a sufficiently large absolute constant $C>0$, such that for any $\delta\in(0,1)$ satisfying
$\frac{\log((n+d)/\delta)}{\sqrt{d}}\left(\alpha+\beta\sqrt{T}+1\right)\leq\frac{1}{C}~,$
it holds with probability at least $1-5nd\delta-\exp(-d/48)$ that $A\in[-1,+1]^{n\times d}$.